OpenStudy (anonymous):

Which of the following shows the equation log2 8 = 3 written in exponential form?

OpenStudy (paxpolaris):

\[\large \log_\color{blue}a x = \color{green}m \]\[\large \implies x = \color{blue}a^\color{green}m\]

OpenStudy (anonymous):

so would it be 3^2 = 8 ???? @PaxPolaris

OpenStudy (paxpolaris):

2^3...

OpenStudy (paxpolaris):

\[8=2^3\] which you can check ...\[2^3=2\times2\times2\]

OpenStudy (anonymous):

2^3 = 8 ? lol @PaxPolaris

OpenStudy (paxpolaris):

right

OpenStudy (anonymous):

thank yaa :)) @PaxPolaris

OpenStudy (anonymous):

so how would u find equation log4 64 = 3 written in exponential form?? @Para_phrenic88